Assume that Bay King Products sells three varieties of canned seafood with the following prices costs: Selling Price per Case $125 $150 $100 Variable Cost per Case $95 $100 $80 Fixed Cost per Month Premium Royal Deluxe Entire firm $246,500 The sales mix (in cases) is 30% Premium, 2090 Royal, and 50% Deluxe.

I just need number 9 & 10, thanks.
Answer #1

1. Weighted Average Contribution Margin=290/10= $29(see note below)

NOTE: Weighted Average Contribution Margin

Particulars

Premium

Royal

Deluxe

Total

Selling price per case

$ 125

$ 150

$ 100

Less: Variable cost per case

$ 95

$ 100

$ 80

Contribution per case

$ 30

$ 50

$ 20

Sales Mix

3

2

5

10

Contribution Margin

$90

$100

$100

$290

Weighted Average Contribution Margin=290/10= $29

  1. Break-Even Point(units)=8500 units(see note below)

NOTE:

At the break even point the Total Contribution=Total Fixed Cost

Assume that units sold to break even are as follows:

Premium=3x , Royal=2x and Deluxe=5x

Therefore Total Contribution=90x+100x+40x=230x

At Break Even Point 290x=$246500x`

i.e x=850

Number of units of Premium= 2550 units

Number of units of Royal= 1700 units

Number of units of Deluxe= 4250 units

  1. Break –Even Point (Sales revenues)=$ 998750(see note below)

Note:

Premium=2550 units*$125=$318750

Royal=1700 units*$150=$2,55,000

Deluxe=4250 units*$100=$4,25,000

  1. Required number of units to earn $20000 profit=9190 units

Note:

Required Profit=$20000

Particulars

Premium

Royal

Deluxe

Total

Selling price per case

$ 125

$ 150

$ 100

Less: Variable cost per case

$ 95

$ 100

$ 80

Contribution per case

$ 30

$ 50

$ 20

Number of units

3x

2x

5x

10x

Total Contribution

$ 90x

$ 100x

$100x

$ 2,66,500

Less: Fixed Costs

$ 2,46,500

Profit

$   20,000

Since $290x=$ 266500

Number of units of premium=2757 units

Number of units of Royal =1838 units

Number of units of Deluxe= 4595 units

Total number of units=9190 units